### v1

@@ 3 Fast-slow analysis of the forced FHN model
The bifurcation diagrams from Figures 1-2 illustrate how slowly varying rhythms could organize spikes in neuronal bursting. The frequency of the input, $\omega$, is independent from the internal properties of the neuron, e.g., from parameters $a$, $b$ and $\varepsilon$. However, since $\omega$ itself is small, it could be defined as the product $\omega=\varepsilon\delta$ with $\varepsilon$ and $\delta$ taking values independently from each other. In this paper we adopt this (re)definition of $\omega$.
@@
Then, in Section 3.4, we fix $\varepsilon>0$ and take the limit $\delta\rightarrow 0$. This limit establishes a two-dimensional fast problem that admits a curve of equilibrium points. Changing to the $(\varepsilon\delta)$-slow-time gives a one-dimensional reduced problem that characterizes the system’s dynamics near hyperbolic points on this curve. In this case, GSPT breaks down at non-hyperbolic points and induces delayed-Hopf dynamics, which contribute to a spike-adding delayed-Hopf-based mechanism that interacts with the aforementioned canard-induced mechanism.
Remark 3.1.
@@ 3.1 Deriving the three-dimensional FHN and the desingularized system
(2a) $\dfrac{dx}{dt}$ $=x-\frac{1}{3}x^{3}-Y$
(2b) $\dfrac{dY}{dt}$ $=\varepsilon(x-bY+ab-Eb\sin\theta-E\delta\cos\theta)$
(2c) $\dfrac{d\theta}{dt}$ $=\varepsilon\delta.$
@@
(3a) $\dfrac{du}{dt}$ $=-v+F(u)$
(3b) $\dfrac{dv}{dt}$ $=\varepsilon(u-bv+\mu-R_{\delta}\cos(\theta-\varphi_{\delta}))$
(3c) $\dfrac{d\theta}{dt}$ $=\varepsilon\delta$
@@
(4) $F(u)$ $=$ $u^{2}-\frac{1}{3}u^{3}$
(5) $\mu$ $=$ $b\left(a+2/3\right)-1\approx 0.233>0$
(6) $R_{\delta}$ $=$ $E\sqrt{b^{2}+\delta^{2}}$
@@
(7a) $\cos(\varphi_{\delta})=\frac{\delta}{\sqrt{b^{2}+\delta^{2}}}$
(7b) $\sin(\varphi_{\delta})=\frac{b}{\sqrt{b^{2}+\delta^{2}}}$
(7c) $\varphi_{\delta}\in\left(0,\frac{\pi}{2}\right),\quad\varphi_{\delta}\to\frac{\pi}{2}\,\,\mbox{as}\,\,\delta\to 0.$
@@ 3.2 Critical manifold
(8a) $\varepsilon\frac{du}{d\tau_{1}}$ $=-v+F(u)$
(8b) $\frac{dv}{d\tau_{1}}$ $=u-bv+\mu-R_{\delta}\cos(\theta-\varphi_{\delta})$
(8c) $\frac{d\theta}{d\tau_{1}}$ $=\delta.$
@@
(9a) $\dfrac{dv}{d\tau_{1}}$ $=u-bv+\mu-R_{\delta}\cos\left(\theta-\varphi_{\delta}\right)$
(9b) $\dfrac{d\theta}{d\tau_{1}}$ $=\delta,$
which is accompanied by the algebraic equation $0=-v+F(u)$. The algebraic equation recapitulates the critical manifold $\mathcal{M}$ for $\theta\in[0,2\pi)$, and so the dynamics for the slow reduced problem (9) occur along $\mathcal{M}$. Fenichel theory [19] states that the dynamics on $\mathcal{M}$ in the reduced slow system persist as slow dynamics for the full system (3), with $0<\varepsilon\ll 1$, along $\mathcal{M}^{\varepsilon}$, a smooth perturbation of $\mathcal{M}$. The stability properties of $\mathcal{M}^{\varepsilon}$ are inherited from hyperbolic regions of $\mathcal{M}$. To distinguish the perturbation $\mathcal{M}^{\varepsilon}$ from the critical manifold $\mathcal{M}$ derived from the singular limit, we will refer to $\mathcal{M}^{\varepsilon}$ as the slow manifold. The distinction of the critical manifold $\mathcal{M}$ from the perturbation $\mathcal{M}^{\varepsilon}$ follows the convention described in [26, Chapter 3]. For non-hyperbolic regions of $\mathcal{M}$, Fenichel theory breaks down, possibly allowing for richer dynamics in the original system near these points. For $\mathcal{M}$, the non-hyperbolic points occur along $\mathcal{L}_{-}$ and $\mathcal{L}_{+}$, called fold lines (see dashed lines in Figure 4A), where $F^{\prime}(u)=0$.
@@
(10) $\dot{u}$ $=R_{\delta}\cos(\theta-\varphi_{\delta})-G(u)$
@@
(11) $G(u)=\mu+u-bF(u)=\mu+u-b\left(u^{2}-\frac{1}{3}u^{3}\right)$
@@ 3.3 Fold points and their stability analysis
Proposition 3.2 (Existence of folded equilibria).
Consider $G$, $\mu$ from (11), (5), $R_{\delta}$, $\varphi_{\delta}$ defined by (6), (7), and parameters $a=0.875$, $b=0.8$, $\varepsilon=0.08$ as in Sec. 2. Then:
@@
(12) $E^{*}_{\ell,\delta}=\frac{ab-(1-2b/3)}{\sqrt{b^{2}+\delta^{2}}}<E.$
@@
(13) $u_{S\ell}$ $=$ $0,\quad v_{S\ell}=0,\quad\theta_{S\ell}=\varphi_{\delta}+\arccos\frac{G(0)}{R_{\delta}}$
(14) $u_{N\ell}$ $=$ $0,\quad v_{N\ell}=0,\quad\theta_{N\ell}=\varphi_{\delta}-\arccos\frac{G(0)}{R_{\delta}}+2\pi\quad(\text{mod}\;2\pi).$
@@
(15) $E^{*}_{r,\delta}=E^{*}_{\ell,\delta}+\frac{2-4b/3}{\sqrt{b^{2}+\delta^{2}}}<E.$
@@
(16) $u_{Sr}$ $=$ $2,\quad v_{Sr}=\frac{4}{3},\quad\theta_{Sr}=\varphi_{\delta}-\arccos\frac{G(2)}{R_{\delta}}+2\pi\quad(\text{mod}\;2\pi)$
(17) $u_{Nr}$ $=$ $2,\quad v_{Nr}=\frac{4}{3},\quad\theta_{Nr}=\varphi_{\delta}+\arccos\frac{G(2)}{R_{\delta}}.$
@@
(18) $E^{*}_{\ell,\delta}<E<E^{*}_{r,\delta}\quad\Longleftrightarrow\quad\frac{ab-(1-2b/3)}{\sqrt{b^{2}+\delta^{2}}}<E<\frac{ab+(1-2b/3)}{\sqrt{b^{2}+\delta^{2}}}.$
Proof 3.3.
@@
Figure 3: Bifurcation diagram for the three-dimensional FHN. Parameter regions according to folded equilibria (here $\omega=\varepsilon\delta$ with $\varepsilon=0.08$): The blue regions (I, II, and III) show the primary parameter region considered in this paper for spike-adding. In regions I, II, III the FHN model admits folded equilibria only on $\mathcal{L}_{-}$. The lines separating regions I from II and II from III are $E=E^{*}_{\ell,\delta}$ and $E=E^{**}_{\ell,\delta}$, with $E^{*}_{\ell,\delta}$, $E^{**}_{\ell,\delta}$ given by (12), (19) respectively. The orange regions (IV, V, and VI) admit additional folded equilibria that lie on $\mathcal{L}_{+}$. The lines separating regions III from V, and V from VI, are $E=E^{*}_{r,\delta}$ and $E=E^{**}_{r,\delta}$ with $E^{*}_{r,\delta}$, $E^{**}_{r,\delta}$ given by Eqs. (15) and (20).
Remark 3.4.
According to formula (12), the lower bound $E^{*}_{\ell,\delta}$ for $E$ increases when $\delta$ decreases. Furthermore, its limiting value can be estimated as $\delta\to 0$ and it is $E^{*}_{\ell,0}=a+\frac{2}{3}-\frac{1}{b}\approx 0.2917>0$. In the parameter plane $(\omega,E)$, where $\omega=\varepsilon\delta$, the curve $E=E^{*}_{\ell,\delta}$ defines a folded saddle-node bifurcation on $\mathcal{L}_{-}$ (Figure 3, this curve separates region I from II). In a similar manner, the curve $E=E^{*}_{r,\delta}$ defines a folded saddle-node bifurcation on $\mathcal{L}_{+}$ and $E^{*}_{r,\delta}$ increases as $\delta\to 0$ with limiting value $E^{*}_{r,0}=E^{*}_{\ell,0}+2/b-4/3\approx 1.4583$ (Figure 3, the curve separates regions III and V, and region II from IV). Clearly, for parameter values below this curve (i.e., the blue-colored parameter range in Figure 3) folded equilibria exist only on $\mathcal{L}_{-}$.
Proposition 3.5 (Type and stability of folded equilibria).
Consider the hypotheses from Proposition 3.2. Then:
@@
(19) $E<E_{\ell,\delta}^{**}=\sqrt{\left(E^{*}_{\ell,\delta}\right)^{2}+\frac{1/(8\delta)^{2}}{b^{2}+\delta^{2}}}$
@@
(20) $E<E_{r,\delta}^{**}=\sqrt{\left(E^{*}_{r,\delta}\right)^{2}+\frac{1/(8\delta)^{2}}{b^{2}+\delta^{2}}}$
@@
The trace is clearly negative for any equilibrium. For $\theta_{N\ell}$ and $\theta_{Sr}$, we have that $\sin(\theta^{*}-\varphi_{\delta})=-\sin[\arccos(G(u^{*})/R_{\delta})]<0$; see (14)-(16) with $G(u^{*})>0$. Since $u^{*}=0<1$ on $\mathcal{L}_{-}$ and $u^{*}=2>1$ on $\mathcal{L}_{+}$, the determinant $\det\left(J_{u\theta}\right)$ is positive for $\theta_{N\ell}$ (thus the folded equilibrium is asymptotically stable) and negative for $\theta_{Sr}$ (a saddle). A similar calculation shows that $\sin(\theta^{*}-\varphi_{\delta})=\sin[\arccos(G(u^{*})/R_{\delta})]>0$ for $\theta_{S\ell}$ (where $u^{*}=1$ thus a saddle) and $\theta_{Nr}$ (where $u^{*}=2$ thus an asymptotically stable equilibrium). We note that the equilibrium is a folded node only when $01-4\det\left(J_{u\theta}\right)>$ 0, which at the equilibrium is equivalent to $1/(8\delta)^{2}+G(u^{*})^{2}>R_{\delta}^{2}$. By direct replacement of $u^{*}=0$ on $\mathcal{L}_{-}$ and $u^{*}=2$ on $\mathcal{L}_{+}$, and using the definition of $G$, $\mu$ from (11), (5) we obtain the results in (ii) and (iii).
Remark 3.7.
@@
System (3)’s dynamics on the critical manifold $\mathcal{M}$ (and, by perturbation, on the slow manifold $\mathcal{M}_{\varepsilon}$) are highly influenced by the attractive/repelling manifolds of the folded nodes and folded saddles. Near the fold points on $\mathcal{L}_{-}$ and $\mathcal{L}_{+}$, the directions of attraction (for the nodes) or of attraction and repulsion (for the saddles) are well approximated by the corresponding eigenvectors. We calculate these quantities below.
Proposition 3.8.
Consider the folded saddles and folded nodes (when they exist) from Proposition 3.5, and $\mu$ defined by (5). Then, in the desingularized system (10)
@@
Proof 3.9.
The eigenvalues and eigenvectors of the equilibria of the desingularized system (10) are obtained from direct calculation from the Jacobian of $J_{u\theta}$ (the matrix $J_{u\theta}$ was introduced in the proof of Proposition 3.5).
The folded singularities on $\mathcal{L}_{-}$ are of particular interest in our analysis. One can estimate the relative order of magnitude of their eigenvalues and eigenvectors when $\delta$ becomes very small (e.g., $\delta\ll\varepsilon$). However, this would be an extreme case of parameter choice that we do not actually consider in this paper (see Figure 2 and Figure 5; we are mostly interested in the cases when $\delta=\mathcal{O}(\sqrt{\varepsilon})$ or $\mathcal{O}(\varepsilon)$). For comparison purposes, we include here a short discussion of the case $\delta\ll\varepsilon$ then comment on the applicability of those results, under certain circumstances, to the case $\delta=\mathcal{O}(\varepsilon)$. By expansion of the eigenvalues in Proposition 3.8, we find that:
@@
Figure 4: Bursting solution with critical and supercritical manifold. A The critical manifold, $\mathcal{M}$, is separated into attracting ($\mathcal{M}_{a-}$, $\mathcal{M}_{a+}$; blue) and repelling ($\mathcal{M}_{r}$; red) regions. The attracting and repelling regions intersect along fold lines, $\mathcal{L}_{\pm}$. The folded node ($N\ell$) and saddle ($S\ell$) are shown as a square and circle, respectively, along the fold lines. The supercritical manifold, $\mathcal{Z}$ (magenta), is a curve along the critical manifold. A bursting trajectory of system (1) with three spikes (forcing parameters $E=0.55$, $\omega=0.0149354$) is shown with $\theta=\omega t$ over one input period. This trajectory contains canard segments, following passage near the folded node and folded saddle, that remain near $\mathcal{M}_{a+}$ for non-negligible time on the fast timescale. B The same trajectory projected onto the $\theta$$x$ axis. Colors and symbols follow the same conventions as in panel A.

### v2

@@ 3 Fast-slow analysis of the forced FHN model
The spike-adding diagrams from Figures 1-2 illustrate how slowly varying rhythms could organize spikes in neuronal bursting. The frequency of the input, $\omega$, is independent from the internal properties of the neuron, e.g., from parameters $a$, $b$ and $\varepsilon$. However, since $\omega$ itself is small, it could be defined as the product $\omega=\varepsilon\delta$ with $\varepsilon$ and $\delta$ taking values independently from each other. In this paper we adopt this (re)definition of $\omega$.
@@
Then, in Section 3.4, we fix $\varepsilon>0$ and take the limit $\delta\rightarrow 0$. This limit establishes a two-dimensional fast problem that admits a curve of equilibrium points. Changing to the $(\varepsilon\delta)$-slow-time gives a one-dimensional reduced problem that characterizes the system’s dynamics near hyperbolic points on this curve. In this case, GSPT breaks down at non-hyperbolic points and induces delayed-Hopf dynamics. We discuss this case for completion purposes. However, we note that a spike-adding delayed-Hopf-based mechanism that, in principle, may interact with the aforementioned canard-induced mechanism is not discernible in our parameter regime of interest, illustrated in Figure 2.
Remark 1
@@ 3.1 Deriving the three-dimensional FHN and the desingularized system
$\dfrac{dx}{dt}$ $=x-\frac{1}{3}x^{3}-Y$ (2a)
$\dfrac{dY}{dt}$ $=\varepsilon(x-bY+ab-Eb\sin\theta-E\delta\cos\theta)$ (2b)
$\dfrac{d\theta}{dt}$ $=\varepsilon\delta.$ (2c)
@@
$\dfrac{du}{dt}$ $=-v+F(u)$ (3a)
$\dfrac{dv}{dt}$ $=\varepsilon(u-bv+\mu-R_{\delta}\cos(\theta-\varphi_{\delta}))$ (3b)
$\dfrac{d\theta}{dt}$ $=\varepsilon\delta$ (3c)
@@
$F(u)$ $=$ $u^{2}-\frac{1}{3}u^{3}$ (4)
$\mu$ $=$ $b\left(a+2/3\right)-1\approx 0.233>0$ (5)
$R_{\delta}$ $=$ $E\sqrt{b^{2}+\delta^{2}}$ (6)
@@
$\cos(\varphi_{\delta})=\frac{\delta}{\sqrt{b^{2}+\delta^{2}}}$ (7a)
$\sin(\varphi_{\delta})=\frac{b}{\sqrt{b^{2}+\delta^{2}}}$ (7b)
$\varphi_{\delta}\in\left(0,\frac{\pi}{2}\right),\quad\varphi_{\delta}\to\frac{\pi}{2}\,\,\mbox{as}\,\,\delta\to 0.$ (7c)
@@ 3.2 Critical manifold
$\varepsilon\frac{du}{d\tau_{1}}$ $=-v+F(u)$ (8a)
$\frac{dv}{d\tau_{1}}$ $=u-bv+\mu-R_{\delta}\cos(\theta-\varphi_{\delta})$ (8b)
$\frac{d\theta}{d\tau_{1}}$ $=\delta.$ (8c)
@@
$\dfrac{dv}{d\tau_{1}}$ $=u-bv+\mu-R_{\delta}\cos\left(\theta-\varphi_{\delta}\right)$ (9a)
$\dfrac{d\theta}{d\tau_{1}}$ $=\delta,$ (9b)
which is accompanied by the algebraic equation $0=-v+F(u)$. The algebraic equation recapitulates the critical manifold $\mathcal{M}$ for $\theta\in[0,2\pi)$, and so the dynamics for the slow reduced problem (9) occur along $\mathcal{M}$. Fenichel theory [30] states that the dynamics on $\mathcal{M}$ in the reduced slow system persist as slow dynamics for the full system (3), with $0<\varepsilon\ll 1$, along $\mathcal{M}^{\varepsilon}$, a smooth perturbation of $\mathcal{M}$. The stability properties of $\mathcal{M}^{\varepsilon}$ are inherited from hyperbolic regions of $\mathcal{M}$. To distinguish the perturbation $\mathcal{M}^{\varepsilon}$ from the critical manifold $\mathcal{M}$ derived from the singular limit, we will refer to $\mathcal{M}^{\varepsilon}$ as the slow manifold. The distinction of the critical manifold $\mathcal{M}$ from the perturbation $\mathcal{M}^{\varepsilon}$ follows the convention described in [32, Chapter 3]. For non-hyperbolic regions of $\mathcal{M}$, Fenichel theory breaks down, possibly allowing for richer dynamics in the original system near these points. For $\mathcal{M}$, the non-hyperbolic points occur along $\mathcal{L}_{-}$ and $\mathcal{L}_{+}$, called fold lines (see dashed lines in Figure 4A), where $F^{\prime}(u)=0$.
@@
$\dot{u}$ $=R_{\delta}\cos(\theta-\varphi_{\delta})-G(u)$ (10)
@@
$G(u)=\mu+u-bF(u)=\mu+u-b\left(u^{2}-\frac{1}{3}u^{3}\right)$ (11)
@@ 3.3 Fold points and their stability analysis
Proposition 1 (Existence of folded equilibria)
Consider $G$, $\mu$ from (11), (5), $R_{\delta}$, $\varphi_{\delta}$ defined by (6), (7), and parameters $a=0.875$, $b=0.8$ as in Sec. 2. Then:
@@
$E^{*}_{\ell,\delta}=\frac{ab-(1-2b/3)}{\sqrt{b^{2}+\delta^{2}}}<E.$ (12)
@@
$u_{S\ell}$ $=$ $0,\quad v_{S\ell}=0,\quad\theta_{S\ell}=\varphi_{\delta}+\arccos\frac{G(0)}{R_{\delta}}$ (13)
$u_{N\ell}$ $=$ $0,\quad v_{N\ell}=0,\quad\theta_{N\ell}=\varphi_{\delta}-\arccos\frac{G(0)}{R_{\delta}}+2\pi\quad(\text{mod}\;2\pi).$ (14)
@@
$E^{*}_{r,\delta}=E^{*}_{\ell,\delta}+\frac{2-4b/3}{\sqrt{b^{2}+\delta^{2}}}<E.$ (15)
@@
$u_{Sr}$ $=$ $2,\quad v_{Sr}=\frac{4}{3},\quad\theta_{Sr}=\varphi_{\delta}-\arccos\frac{G(2)}{R_{\delta}}+2\pi\quad(\text{mod}\;2\pi)$ (16)
$u_{Nr}$ $=$ $2,\quad v_{Nr}=\frac{4}{3},\quad\theta_{Nr}=\varphi_{\delta}+\arccos\frac{G(2)}{R_{\delta}}.$ (17)
@@
$E^{*}_{\ell,\delta}<E<E^{*}_{r,\delta}\quad\Longleftrightarrow\quad\frac{ab-(1-2b/3)}{\sqrt{b^{2}+\delta^{2}}}<E<\frac{ab+(1-2b/3)}{\sqrt{b^{2}+\delta^{2}}}.$ (18)
Proof 1
@@
Figure 3: Bifurcation diagram for the three-dimensional FHN. Parameter regions according to folded equilibria: The blue regions (I, II, and III) show the primary parameter region considered in this paper for spike-adding. In regions I, II, III the FHN model admits folded equilibria only on $\mathcal{L}_{-}$. The lines separating regions I from II and II from III are $E=E^{*}_{\ell,\delta}$ and $E=E^{**}_{\ell,\delta}$, with $E^{*}_{\ell,\delta}$, $E^{**}_{\ell,\delta}$ given by (12), (19) respectively. The orange regions (IV, V, and VI) admit additional folded equilibria that lie on $\mathcal{L}_{+}$. The lines separating regions III from V, and V from VI, are $E=E^{*}_{r,\delta}$ and $E=E^{**}_{r,\delta}$ with $E^{*}_{r,\delta}$, $E^{**}_{r,\delta}$ given by Eqs. (15) and (20). Note that the horizontal axis closely resembles that of Figure 2 where $\omega=\varepsilon\delta\in[0,0.08]$ and $\varepsilon=0.08$.
Remark 2
[…]
Proposition 2 (Type and stability of folded equilibria)
Consider the hypotheses from Proposition 1. Then:
@@
$E<E_{\ell,\delta}^{**}=\sqrt{\left(E^{*}_{\ell,\delta}\right)^{2}+\frac{1/(8\delta)^{2}}{b^{2}+\delta^{2}}}$ (19)
@@
$E<E_{r,\delta}^{**}=\sqrt{\left(E^{*}_{r,\delta}\right)^{2}+\frac{1/(8\delta)^{2}}{b^{2}+\delta^{2}}}$ (20)
@@
The trace is clearly negative for any equilibrium. For $\theta_{N\ell}$ and $\theta_{Sr}$, we have that $\sin(\theta^{*}-\varphi_{\delta})=-\sin[\arccos(G(u^{*})/R_{\delta})]<0$; see (14)-(16) with $G(u^{*})>0$. Since $u^{*}=0<1$ on $\mathcal{L}_{-}$ and $u^{*}=2>1$ on $\mathcal{L}_{+}$, the determinant $\det\left(J_{u\theta}\right)$ is positive for $\theta_{N\ell}$ (thus the folded equilibrium is asymptotically stable) and negative for $\theta_{Sr}$ (a folded saddle). A similar calculation shows that $\sin(\theta^{*}-\varphi_{\delta})=\sin[\arccos(G(u^{*})/R_{\delta})]>0$ for $\theta_{S\ell}$ (where $u^{*}=1$ thus a folded saddle) and $\theta_{Nr}$ (where $u^{*}=2$ thus an asymptotically stable equilibrium). We note that the equilibrium is a folded node only when $01-4\det\left(J_{u\theta}\right)>$ 0, which at the equilibrium is equivalent to $1/(8\delta)^{2}+G(u^{*})^{2}>R_{\delta}^{2}$. By direct replacement of $u^{*}=0$ on $\mathcal{L}_{-}$ and $u^{*}=2$ on $\mathcal{L}_{+}$, and using the definition of $G$, $\mu$ from (11), (5) we obtain the results in (ii) and (iii).
Remark 3
@@
System (3)’s dynamics on the critical manifold $\mathcal{M}$ (and, by perturbation, on the slow manifold $\mathcal{M}_{\varepsilon}$) are highly influenced by the attractive/repelling manifolds of the folded nodes and folded saddles. Near the fold points on $\mathcal{L}_{-}$ and $\mathcal{L}_{+}$, the directions of attraction (for the folded nodes) or of attraction and repulsion (for the folded saddles) are well approximated by the corresponding eigenvectors. We calculate these quantities below.
Proposition 3
Consider the folded saddles and folded nodes (when they exist) from Proposition 2, and $\mu$ defined by (5). Then, in the desingularized system (10)
@@
Proof 3
The eigenvalues and eigenvectors of the equilibria of the desingularized system (10) are obtained from direct calculation from the Jacobian of $J_{u\theta}$ (the matrix $J_{u\theta}$ was introduced in the proof of Proposition 2).
The folded singularities on $\mathcal{L}_{-}$ are of particular interest in our analysis. One can estimate the relative order of magnitude of their eigenvalues and eigenvectors when $\delta$ becomes very small (e.g., $\delta\ll\varepsilon$). However, this would be an extreme case of parameter choice that we do not actually consider in this paper (see Figure 2 and Figure 5; we are mostly interested in the cases when $\delta=\mathcal{O}(\sqrt{\varepsilon})$ or $\mathcal{O}(\varepsilon)$). For comparison purposes, we include here a short discussion of the case $\delta\ll\varepsilon$ then comment on the applicability of those results, under certain circumstances, to the case $\delta=\mathcal{O}(\varepsilon)$. By expansion of the eigenvalues in Proposition 3, we find that: